OpenStudy (anonymous):

A school record in a certain race in 1960 was 3.8 minutes. In 1980 the school record in the same event was 3.65 minutes. choose the linear equation for these data points: A) y=.0075x -18.5 B) y=-.0075x + 18.5 C) y=.0075x+18.5

OpenStudy (anonymous):

help me ..................

OpenStudy (anonymous):

times it by 2

OpenStudy (anonymous):

how about y'all just shut up if you're not going to help..

OpenStudy (anonymous):

me

OpenStudy (anonymous):

too

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

make me

OpenStudy (anonymous):

alright ni realllllyy need help and she will cut you

OpenStudy (anonymous):

and how ols are yall acting like 10 yr oldsss and its not wsh its wish ...

OpenStudy (anonymous):

look, bro. you don't know me and i'm not about to fight a CHILD on the internet. delaine, i'm not going to cut anyone. "poopster09", please leave us alone. because you're just downright obnoxious.

OpenStudy (anonymous):

hahahahaahahahahhahahahhahaha

OpenStudy (anonymous):

it would be b. from 1960 to 1980 the time decrease by 0.15 minutes, so slope would be \[ -(0.15/20)=-0.0075\] as b is the only equation with negative slope, it has to be the right one. Also you could just insert the numbers (x=year) if x=1960 then: \[-0.0075*1960 +18.5 = -14.7 + 18.5 = 3.8\] if x=1980 then \[-0.0075*1980 +18.5 = -14.85 + 18.5 = 3.65\] so you have double proof of b being right ;)

OpenStudy (anonymous):

ahaahahah

OpenStudy (anonymous):

cool story, bro.